# Networked Communication for Decentralised Cooperative Agents in Mean-Field Control

### v1

@@ E.3 Additional experiments and ablations
Figure 3: All communication links suffer a 90% probability of failure, including in the centralised case, where the link between the central learner and the rest of the population may fail. $C_{e}=C_{r}=C_{p}=1$. The centralised population, which in the standard setting matched networked performance only in the ‘cluster’ game, now learns slower even in this game, due to suffering from the single point of failure. Our networked scheme appears robust to the failures in all games, with only small differences to performance in the standard setting. In fact, several broadcast radii perform better in the ‘shape formation’ game with these failures than without, probably because they permit greater diversity policies while still having an advantage over purely independent learners (as discussed in Sec. 5.2). However, the smallest broadcast radius (green, 0.2) does drop in performance in this game, which might be expected given it now acts similarly to the independent case. Networked populations appear to have less variance in this setting than in the standard setting, at least in the first four games. This is likely because the communication failures prevent both particularly high and particularly low performing policies from spreading fast in the population, preventing large performance fluctuations and smoothing learning progress. Meanwhile a centralised population still has large variance even with communication failures, due to enforcing the adoption of an arbitrarily-chosen consensus policy - in some games variance is higher in this setting (though in some it may be marginally lower). This points to an additional benefit of our networked scheme over the centralised case.
@@
Figure 4: Standard algorithms but $C_{e}=C_{r}=C_{p}=$ 10. As is expected, in the coordination games the networked agents with lower broadcast radii now receive returns almost as high as those with larger radii, albeit at the cost of greater variance (as there may be some noise in the quality of the policy that gets spread to the whole population as a result of more communication rounds). In the ‘target selection’ game, now all networked populations outperform the centralised agents. In the anti-coordination ‘target coverage’ game, the smaller broadcast radii (green, 0.2; red, 0.4; purple, 0.6) receive slightly lower returns than before, since the additional communication rounds now make policy alignment more likely, reducing $f_{d}$ as per Def 10. The same is true of the smallest radius population (green, 0.2) in the ‘shape formation’ game, which receives a lower return than before. Nevertheless, all networked populations receive higher returns than the independent agents in all games, and also than the centralised population in all but the ‘cluster’ game. This shows that in our experimental settings there is a very large benefit to a single communication round, with limited benefit to increasing the algorithms’ time complexity with additional communication rounds.
@@
Figure 5: Standard algorithms but $C_{e}=C_{r}=C_{p}=$ 50. Having 50 communication rounds does not appear to significantly change networked performance compared to 10 rounds (Fig. 4), with most increases or decreases in average return appearing within the margin of error. Most notably, the largest broadcast radius (pink, 1.0) receives slightly lower return now than with 10 rounds in the ‘disperse’ game, while pink (1.0), brown (0.8) and green (0.2) receive lower returns and have higher variance now in the ‘beach bar’ game. As in the case of $C_{e}=C_{r}=C_{p}=10$, additional communication rounds make policy alignment more likely, reducing $f_{d}$ as per Def 10. Nevertheless, all networked populations receive higher returns than the independent agents in all games, and also than the centralised population in all but the ‘cluster’ game. This shows that in our experimental settings there is a very large benefit to a single communication round, with limited benefit to increasing the algorithms’ time complexity with additional communication rounds.
@@
Figure 6: Ablation study on population-independent policies. No agents, including centralised and networked ones, observe the empirical mean field, and all receive a vector of zeros in its place (so as to keep the neural networks the same size as in the standard setting). $C_{r}=C_{p}=1$. In our stationary games, networked populations do not appear to perform substantially differently to the standard population-dependent setting, though some radii (red, 0.4; pink, 1.0) appear to perform slightly better in the ‘shape formation’ game. On the other hand, in the coordination games, and particularly the ‘target selection’ game, the centralised population receives a significantly lower return in this setting.
@@
Figure 7: Ablation study of Alg. 3 for estimating the empirical mean field - all agents, including independent ones, directly receive the true global empirical mean field. $C_{r}=C_{p}=1$. This does not appear to change performance in the networked populations (apart from greater variance here in the ‘shape formation’ game), nor does it help independent agents. This may be evidence that Alg. 3 enables networked agents to accurately estimate the global mean field from local observations. However, our ablation study on population-independent policies (Fig 6) suggests that not observing the mean field does not markedly disadvantage agents in our experimental settings in any case (apart from for the centralised populations in the coordination games). Therefore further evidence is required in settings that require population-dependent policies to confirm the efficacy of Alg. 3.
@@
Figure 8: Ablation study for observation of true/estimated global average reward $\hat{r}$/$\tilde{\hat{r}}^{i}_{t}$, where all agents, including centralised ones, only have access to $r^{i}_{t}$, where in the centralised case $i=1$. $C_{e}=C_{p}=1$. The greatest effect of this is on the centralised (blue) case, which performs much worse in the ‘target selection’ game, and with higher variance in the ‘cluster’ and ‘beach bar’ games. The networked agents appear more robust, though do experience a slight performance decrease, mostly among populations with the largest broadcast radii (pink, 1.0; brown, 0.8), i.e. those most similar to the centralised case in terms of $\tilde{\hat{r}}^{i}_{t}$, as might be expected. In particular, note the greater variance of pink (1.0) in the ‘target selection’ game; slower learning and higher variance of pink (1.0) and brown (0.8) in the ‘beach bar’ game; lower returns for pink (1.0) and brown (0.8) in the ‘shape formation’ game; and slower learning and convergence of the smallest radii (green, 0.2; red, 0.4) in the ‘target coverage’ game. This all demonstrates the usefulness and efficacy of our novel Alg. 1.
@@
Figure 9: Ablation study for Alg. 1 for estimating the true global average reward. All agents, including both networked and independent ones, directly receive the true global average reward such that $\tilde{\hat{r}}^{i}_{t}=\hat{r}$. Access to the true average reward does not help networked (or independent) agents to improve their returns, demonstrating that our novel Alg. 1 already affords networked populations robustness against the lack of access to this global information (having this global information would be an unrealistic assumption in practice). In fact, networked populations’ performance actually seems to be worse with this global information in the ‘shape formation’ game, particularly in the case with the smallest broadcast radius (green, 0.2), but perhaps not by a statistically significant amount.
@@
Figure 10: Ablation study of the choice of $\tau^{comm}_{k}$. Here $\forall k\;\tau^{comm}_{k}=$ 1e-18 (i.e. $\tau^{comm}_{k}\rightarrow 0$), rather than linearly increasing from 0.001 to 1 across the $K$ iterations as in all other experiments (see Table 1). $C_{e}=C_{r}=C_{p}=1$. In this setting, networked agents continue to outperform the centralised (blue) and independent (orange) populations in all games (except the ‘cluster game’), but otherwise generally appear to receive lower average returns and with greater variance. This is because Ass. 2 on the quality of the finite-step approximations $\sigma^{i}_{k+1}$}${}_{i=1}^{N}$ = $\{\hat{V}^{i}(\boldsymbol{\pi}_{k+1},\mu_{0};E)\}_{i=1}^{N}$ does not always apply in practice, meaning the policy estimated to perform the best may not actually be a good update, such that enforcing the adoption of this policy can lead to noisy, unstable learning. Using a higher temperature value smooths out this noise. Moreover, using $\tau^{comm}_{k}\rightarrow 0$ effectively enforces consensus on a single policy for the finite population in the networked case, which in anti-coordination games may also reduce the average return. This all provides empirical evidence for our scheme for $\tau^{comm}_{k}$, but further optimising the choice might lead to additional performance increase.

### v2

@@ 6.3 Results and discussion
Figure 3: All communication links suffer a 90% probability of failure, including in the central-agent case, where the link between the central learner and the rest of the population may fail. $C_{e}=C_{r}=C_{p}=1$. The central-agent population, which in the standard setting matched networked performance only in the ‘cluster’ game, now learns slower even in this game, due to suffering from the single point of failure. Our networked scheme appears robust to the failures in all games, with only small differences compared to performance in the standard setting. In fact, several broadcast radii appear to perform better in the ‘shape formation’ game with these failures than without (though not significantly so), probably because they permit greater diversity policies while still having an advantage over purely independent learners (as discussed in the body of Sec. 6.3). However, the smallest broadcast radius (green, 0.2) does drop in performance in this game, which might be expected given it now acts similarly to the independent case. Networked populations appear to have less variance in this setting than in the standard setting, at least in the first four games. This is possibly because the communication failures prevent both particularly high and particularly low performing policies from spreading fast in the population, preventing large performance fluctuations and smoothing learning progress. Meanwhile a central-agent population still has large variance even with communication failures, due to enforcing the adoption of an arbitrarily-chosen consensus policy - in some games variance is higher in this setting (though in some it may be marginally lower). This points to an additional benefit of our networked scheme over the central-agent case.
@@
Figure 4: Standard algorithms but $C_{e}=C_{r}=C_{p}=$ 10. As is expected, in the coordination games the networked agents with lower broadcast radii now receive returns almost as high as those with larger radii, albeit at the cost of greater variance (having more communication rounds leads to greater policy consensus in the population at each iteration of the outer loop, and there may be some noise in the quality of these consensus policies). In the ‘target selection’ game, now all networked populations appear to outperform the central-agent (orange) population, though again with high variance. In the anti-coordination ‘target coverage’ game, the smaller broadcast radii (green, 0.2; red, 0.4; purple, 0.6) receive slightly lower returns than before, since the additional communication rounds now make policy alignment more likely, reducing $f_{d}$ as per Def 5.3. The same is true of the smallest radius population (green, 0.2) in the ‘shape formation’ game, which receives a lower return than before. This reflects the discussion in Sec. 6.3 regarding the detrimental effect of additional policy adoption once the maximum base return has been achieved in anti-coordination games. Nevertheless, all networked populations receive higher returns than the independent agents in all games, and also than the central-agent population in all but the ‘cluster’ game. This shows that in our experimental settings there is a very large benefit to a single communication round, with limited benefit to increasing the algorithms’ time complexity with additional communication rounds.
@@
Figure 5: Standard algorithms but $C_{e}=C_{r}=C_{p}=$ 50. Having 50 communication rounds does not appear to significantly change networked performance compared to 10 rounds (Fig. 4), with most increases or decreases in average return appearing within the margin of error. Most notably, the largest broadcast radius (pink, 1.0) receives slightly lower return now than with 10 rounds in the ‘disperse’ game, while pink (1.0), brown (0.8) and green (0.2) receive lower returns and have higher variance now in the ‘beach bar’ game. As in the case of $C_{e}=C_{r}=C_{p}=10$, additional communication rounds make policy alignment more likely, reducing $f_{d}$ as per Def 5.3. This reflects the discussion in Sec. 6.3 regarding the detrimental effect of additional policy adoption once the maximum base return has been achieved in anti-coordination games. Nevertheless, all networked populations receive higher returns than the independent agents in all games, and also than the central-agent population in all but the ‘cluster’ game. This shows that in our experimental settings there is a very large benefit to a single communication round, with limited benefit to increasing the algorithms’ time complexity with additional communication rounds.
@@
Figure 6: Ablation study on population-independent policies. No agents, including centralised and networked ones, observe the empirical mean field, and all receive a vector of zeros in its place (so as to keep the neural networks the same size as in the standard setting). $C_{r}=C_{p}=1$. Networked populations do not appear to perform substantially differently to the standard population-dependent setting, though some radii (red, 0.4; pink, 1.0) appear to perform slightly better in the ‘shape formation’ game. This is likely because all of our games have stationary solutions, such that observing the mean field is not actually necessary, even if it could potentially be useful (see Sec. 3.1 for discussion of the conception of MFC as a central planner trying to guide the population to a distribution that maximises the expected return). Indeed, in the coordination games, and particularly the ‘target selection’ game, the central-agent population receives a lower return in this setting, whereas our networked populations are robust to this change.
@@
Figure 7: Ablation study of Alg. 4 for estimating the empirical mean field - all agents, including independent ones, directly receive the true global empirical mean field. $C_{r}=C_{p}=1$. This does not appear to change performance in the networked populations (apart from greater variance here in the ‘shape formation’ game), nor does it help independent agents. This may be evidence that Alg. 4 enables networked agents to accurately estimate the global mean field from local observations. However, our ablation study on population-independent policies (Fig. 6) suggests that not observing the mean field does not markedly disadvantage agents in our experimental settings in any case (apart from for the central-agent populations in the coordination games). This is likely because all of our games have stationary solutions, such that observing the mean field is not necessary. Therefore, in order to confirm the efficacy of Alg. 4 for estimating the mean field, further evidence is perhaps needed in MFC settings that require population-dependent policies, though benjamin2024networkedapproximation has already confirmed this for non-stationary games in the non-cooperative MFG setting.
@@
Figure 8: Ablation study for observation of true/estimated global average reward $\hat{r}$$\tilde{\hat{r}}^{i}_{t}$, where all agents, including centralised ones, only have access to $r^{i}_{t}$, where in the central-agent case $i=1$. $C_{e}=C_{p}=1$. The greatest effect of this is on the central-agent (blue) populations, which perform much worse in the ‘target selection’ game, and with higher variance in the ‘cluster’ and ‘beach bar’ games, i.e. they suffer without access to the global average reward. The networked agents appear more robust to the loss of the (estimated) average reward, pointing to an additional benefit of the policy communication scheme, though do experience a slight performance decrease, mostly among populations with the largest broadcast radii (pink, 1.0; brown, 0.8), i.e. those most similar to the central-agent case in terms of $\tilde{\hat{r}}^{i}_{t}$, as might be expected. In particular, note the greater variance of pink (1.0) in the ‘target selection’ game; slower learning and higher variance of pink (1.0) and brown (0.8) in the ‘beach bar’ game; lower returns for pink (1.0) and brown (0.8) in the ‘shape formation’ game; and slower learning and convergence of the smallest radii (green, 0.2; red, 0.4) in the ‘target coverage’ game. This all demonstrates the usefulness and efficacy of our novel Alg. 1 for decentralised estimation of the global average reward.
@@
Figure 9: Ablation study for Alg. 1 for estimating the true global average reward. All agents, including both networked and independent ones, directly receive the true global average reward such that $\tilde{\hat{r}}^{i}_{t}=\hat{r}$. $C_{e}=C_{p}=1$. Access to the true average reward does not help networked agents to improve their returns, demonstrating that our novel Alg. 1 already affords networked populations robustness against the lack of access to this global information (having this global information would be an unrealistic assumption in practice). Access to the true average reward also does not help independent agents to improve their returns, suggesting the policy communication scheme is the dominant factor in improving the performance of decentralised agents.
@@
Figure 10: Ablation study of the choice of $\tau^{comm}_{k}$. Here $\forall k\;\tau^{comm}_{k}=$ 1e-18 (i.e. $\tau^{comm}_{k}$ is close to 0, turning the softmax into a max function), rather than linearly increasing from 0.001 to 1 across the $K$ iterations as in all other experiments (see Table 1). $C_{e}=C_{r}=C_{p}=1$. In this setting, networked agents continue to outperform the central-agent (blue) and independent (orange) populations in all games except the ‘cluster’ game, but otherwise generally appear to receive lower average returns and with greater variance. This is because Assumption 5.6 on the quality of the finite-step approximations $\sigma^{i}_{k+1}$$_{i=1}^{N}$ = $\{\hat{V}^{i}(\bm{\pi}_{k+1},\mu_{t};E)\}_{i=1}^{N}$ may not always apply in practice, especially as policies get more deterministic closer to convergence. This means the policy estimated to perform the best may not actually be among the best updates, such that enforcing the adoption of this policy can lead to noisy, unstable learning. Using a higher temperature value smooths out this noise (while having a lower temperature at the beginning of training ensures faster learning when the difference in the quality of nascent policies is likely to be more stark, hence our inverse annealing scheme). Moreover, using $\tau^{comm}_{k}$ close to 0 more effectively enforces consensus on a single policy in the networked case, which in anti-coordination games may also reduce the average return (see Sec. 6.3). This all provides empirical evidence for our scheme for $\tau^{comm}_{k}$, but further optimising the choice might lead to additional performance increase.
7 Conclusion
We provided the first algorithms for decentralised training in MFC, as well as the first for online learning in MFC from a single non-episodic run of the empirical system. We did so by modifying existing algorithms for the MFG setting, and contributing a novel algorithm for estimating the global average reward via local communication. We proved theoretically that networked communication accelerates learning over both the independent and central-agent architectures. We supported this with extensive numerical results, accompanied by ablation studies and discussion of the empirical effects of communication radii.
Our work follows the gold standard in MFGs by presenting experiments on grid world toy environments. Nevertheless future work includes experiments in other games, including non-stationary games, more realistic environments and ones where both the transition function and the reward function depend on the mean field. Please note, however, that benjamin2024networkedapproximation already demonstrated in the MFG setting that the communication scheme affords faster learning when both the transition and reward functions depend on the mean field.
In Sec. 5 we give theoretical results showing that our networked algorithm can outperform the central-agent and independent alternative. We leave more general theoretical results, such as proofs of convergence and sample complexity, for future work.
Our algorithms contain numerous inner loops and thus require synchronisation between communicating agents. Our ablation studies of the sub-routines and our experiment on robustness to communication failures (Fig. 3) indicate that synchronisation failure is not necessarily a problem in practice, but future work nevertheless lies in simplifying the nested loops of our algorithms.
In grid-world settings such as those in our experiments, passing the (estimated or true global) mean-field distribution as a flat vector to the Q-network ignores the geometric structure of the problem. perrin2022generalization therefore proposes to create an embedding of the distribution by first passing the vector to a convolutional neural network, essentially treating the categorical distribution as an image. This technique is also followed in wu2024populationaware (for their additional experiments, but not in the main body of their paper). As future work, we can test whether such a method increases the usefulness of observing the mean field in population-dependent policies, and therefore increases the importance of be able to accurately estimate the global mean field via Alg. 4.
Broader Impact Statement
We identified no specific ethical concerns regarding our work, which explores new game theoretical and machine learning algorithms in general settings.